OpenStudy (anonymous):

A printing press produces 125 pages of a book every 3 minutes. At this rate what is the maximum number of complete 300 page books that this press can produce in 4 hours?

OpenStudy (anonymous):

\[\frac{\frac{4*60}{3}125}{300}=\frac{100}{3}=33.3 \text{ books} \]

OpenStudy (jlastino):

If it produces 125 pages for every 3 minutes, then it produces 2500 pages per hour or 10,000 books in 4 hour The problem asks for the maximum numbe of 300 page books divide 10,000 by 300 33.33 Well there can't be 33.33 books So we cut of the .33 It can only print 33 books

OpenStudy (jlastino):

*off

OpenStudy (anonymous):

Sorry, over looked the word "complete' in the problem statement.

OpenStudy (anonymous):

\[\text{Floor}\left[\frac{\frac{4*60}{3}125}{300}\right]=33 \]